# Critical point representation of the mutual information in the sparse stochastic block model

### v1

@@ 1.1 Informal summary of the main results
We denote by $\mathcal{M}_{+}$ the space of non-negative measures over $[-1,1]$ with finite total mass, and by $\mathcal{M}_{p}$ the set of probability measures with mean $\overline{m}=2p-1$,
@@
Our main results will be stated in terms of a fixed point of the functional $\Gamma_{t,\mu}:\mathcal{M}_{p}\to\mathcal{M}_{p}$, which is defined for each $t\geq 0$ and $\mu\in\mathcal{M}_{+}$ according to the following procedure. Fixing $\nu\in\mathcal{M}_{p}$, we sample $\epsilon^{*}$ according to $P^{*}$, and conditionally on $\epsilon^{*}$, we let $\Pi_{t,\mu}(\nu)$ be a Poisson point process with intensity measure ${(c+\Delta\epsilon^{*}x)}\,\mathrm{d}(\mu+t\nu)(x)$; we then set
$\Gamma_{t,\mu}(\nu):=\text{Law}\Bigg{(}\frac{\int_{\Sigma_{1}}\epsilon e^{-\epsilon\overline{m}}\prod_{x\in\Pi_{t,\mu}(\nu)}(c+\Delta\epsilon x)\,\mathrm{d}P^{*}(\epsilon)}{\int_{\Sigma_{1}}e^{-\epsilon\overline{m}}\prod_{x\in\Pi_{t,\mu}(\nu)}(c+\Delta\epsilon x)\,\mathrm{d}P^{*}(\epsilon)}\Bigg{)}.$ (1.6)
@@
    conditionally on $(x_{i})_{i\geq 1}$, generate ±$\pm 1$-valued random variables $\sigma_{i}^{\ell}$ with mean $x_{i}$, independently over $i,\ell\geq 1$.
@@
A complete characterization of the limit of the mutual information has already been achieved in a number of cases. The disassortative case was settled in [2, 15] when $p=1/2$, and in [18] for arbitrary $p$. The technique in [18] is based on partial differential equations (see also [19] for a presentation of the approach), and the key point is that in the disassortative case, the non-linearity appearing in this equation is convex. We expect this complete characterization to extend to settings with more than two communities, as long as the corresponding non-linearity remains convex. This would be analogous to [11, Theorem 1.1] in the context of spin glasses, a result which itself builds upon [9, 22, 42, 43, 44, 45, 46, 47, 52]. Moreover, whenever this convexity property holds, the limit mutual information can be expressed as a variational formula. From the point of view of partial differential equations, this relates to the validity of a Hopf-Lax representation of the solution, see [20, Theorem 1.3].
@@
We expect the method presented here, primarily based on cavity calculations, to be robust to model variations such as situations involving more than two communities, and we also expect the lower bound on the mutual information obtained in [18] to generalize similarly. The results we present here and those in [18] are analogous to those obtained for spin glasses in [11] and in [39, 40] respectively. We have preferred to stick to a two-community model here to facilitate the presentation, but as an illustration of the robustness of the approach, we will at least indicate as we proceed the few small adaptations that need to be implemented in order to cover the case when then measure $P^{*}$ is arbitrary with support in $[-1,1]$ in place of $\{-1,1\}$.
@@
In order to make further progress and obtain a full characterization of the limit mutual information in general, one must first ask what a good candidate for the limit could be. One proposal based on partial differential equations was put forward in [18]. It would of course be convenient to identify a more explicit formulation of the candidate limit. For dense versions of the problem we consider, where the average degree of the graph diverges to infinity, a sort of central limit theorem takes place at the level of each node, and one can relate the problem to a simpler setup such as that of rank-one matrix estimation in the presence of additive Gaussian noise, as shown in [17, 26] (see also [19, Section 4.5]). For the latter class of problems, a saddle-point variational formula was shown to be valid in a very wide range of settings in [14] (see also [19, Chapter 4]), building upon [4, 5, 6, 10, 13, 12, 23, 26, 27, 28, 29, 31, 32, 37, 38, 50, 51]. From the point of view of partial differential equations, this corresponds to the possibility to write a Hopf formula based on the convexity of the initial condition (as opposed to the convexity of the non-linearity for the Hopf-Lax formula, cf. [19, Theorems 3.8 and 3.13]). However, the convexity of the relevant “initial condition” in our setting is invalid [25], and as explained around [18, (1.49)], this implies that this tentative saddle-point Hopf formula is invalid in our setting with sparse graph connectivity. An alternative candidate limit for the mutual information could be considered by extending the Hopf-Lax variational formula known to be valid in the disassortative case. While we could not rule this out in the two-community case, we show here that this candidate variational formula is invalid in general. To do so, we focus on a version of the two-community problem in which the graph additionally possesses a bipartite structure. The model can be thought of as involving four communities, and it is inspired by the bipartite spin-glass model investigated in [39] (see in particular [39, Section 6]). In this context, we show that a naive generalization of the formula valid in the standard disassortative two-community case would lead to a prediction that is in contradiction with the main results of this paper.
@@ 1.2 Precise statement of the main results
We now turn to a more precise description of our main results. As in [18], our starting point is to embed the community-detection problem under consideration into a richer family parametrized by a pair $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$. The conjecture in [18] is that, up to a simple affine transformation, the limit $f(t,\mu)$ of the re-scaled mutual information between the signal and the observation solves an explicit Hamilton-Jacobi equation. As is well-known in finite dimensions, first-order Hamilton-Jacobi equations can be solved for a short time using the method of characteristics (see for instance [19, Exercise 3.10 and solution]). This method allows us to uncover the value of the solution to the equation along each characteristic line separately. At larger times though, difficulties occur because characteristic lines may intersect, and will typically not agree on what the value of the solution ought to be. Our main result can be interpreted as saying that, while this might indeed occur, there is always one characteristic line that prescribes the correct value for $f(t,\mu)$. In the language of [19, Section 3.5], one may say that the graph of $f$ always belongs to the wavefront of the equation.
@@
By this we mean that the law of $\sigma^{*}$ given the observation of $\mathcal{D}_{N}^{t}$ is described by a formula analogous to that in (1.9), with $H_{N}^{\circ}$ replaced by $H_{N}^{t}$ there. By [18, Appendix A], the difference between the free energies $\overline{F}_{N}^{\circ}$ and $\overline{F}_{N}(1)$ converges to zero as $N$ tends to infinity. The main advantage of this construction is that it allows one to study derivatives of the free energy with respect to the “time” parameter $t$. It is also convenient to study derivatives of the free energy with respect to an additional parameter denoted by $\mu\in\mathcal{M}_{+}$. Given $\mu\in\mathcal{M}_{+}$, we let $(\Lambda_{i}(N\mu))_{i\leq N}$ be independent Poisson point processes with intensity measure $N\mu$, and we let the $\mu$-dependent Hamiltonian $H_{N}^{\mu}$ on $\Sigma_{N}$ be given by
@@
For a review of Poisson point processes, we refer the reader to [19, Chapter 5]. The function $H_{N}^{\mu}$ is the Hamiltonian associated with the task of inferring the signal $\sigma^{*}$ from the data
@@
where the randomness in these two datasets is taken to be independent conditionally on $\sigma^{*}$. We recall that by [18, Appendix A], the difference between the free energy $\overline{F}_{N}^{\circ}$ in (1.11) and $\overline{F}_{N}(1,0)$ tends to zero as $N$ tends to infinity. The problem of finding the asymptotic value of the mutual information (1.7) is therefore a specific instance of the task of determining the limit of the enriched free energy (1.22).
@@
for Poisson point processes $\Pi_{\pm}(\mu)$ with respective intensity measures $(c\pm\Delta x)\,\mathrm{d}\mu(x)$ on $[-1,1]$. The function $\psi$ is the initial condition of the partial differential equation that was identified in [18], since it arises as the large-$N$ limit of $\overline{F}_{N}(0,\mu)$. Our first main result states that the limit of the free energy, assuming it exists, is given by the functional (1.26) evaluated at one of its critical points. As mentioned in the previous subsection, it will be possible to recast the critical-point condition as a fixed-point equation for the operator $\Gamma_{t,\mu}$ defined in (1.6), so that our first main result reads as follows.
@@
[…]
Our next result states that for most choices of $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$, we can identify a fixed point of the map $\nu\mapsto\Gamma_{t,\mu}(\nu)$ for which the equality (1.28) holds, and for which we can also identify the limit law of $(\sigma^{\ell})_{\ell\geq 1}$ after a small perturbation of the inference problem. More precisely, we will show that this can be done at any point $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathrm{Reg}(\mathcal{M}_{+})$ at which the limit free energy is Gateaux differentiable; we refer to Section 2 for a definition of Gateaux differentiability and a precise discussion of why the limit free energy is Gateaux differentiable at most points. The convergence in law of $(\sigma^{\ell}_{i})_{i\leq N,\ell\geq 1}$ is in the sense of finite-dimensional distributions.
@@ Acknowledgements.
We would like to warmly thank Dmitry Panchenko and Jean Barbier for sharing their notes [7] on the free energy in the disassortative sparse stochastic block model with us, which helped us with many of the computations in Section 4.
@@ 2 Regularity properties of the free energy
In this section we discuss three regularity properties of the enriched free energy (1.22). In Section 2.1, we establish a Lipschitz continuity bound for the enriched free energy which implies in particular that the free energy restricted to the space of probability measures is Lipschitz continuous with respect to the Wasserstein distance. In Section 2.2, we prove that a Lipschitz continuous function on the space of cumulative distribution functions on $[-1,1]$, viewed as a subset of $L^{2}[-1,1]$, is Gateaux differentiable on a dense subset of its domain. Finally, in Section 2.3, we leverage a convexity property of the free energy to prove that the sequence of derivatives of the free energy converges to the derivative of the limit free energy at any point of differentiability of said limit. We now state the three main results of this section. Since the ideas developed in these proofs will not reappear later, the reader may consider skipping these proofs on first reading.
@@
$\big{\lvert}\overline{F}_{N}(t,\mu)-\overline{F}_{N}(t^{\prime},\mu^{\prime})\big{\rvert}\leq C\big{(}\lvert t-t^{\prime}\rvert+\big{\lvert}\mu[-1,1]-\mu^{\prime}[-1,1]\big{\rvert}+\mu[-1,1]\lvert F_{\overline{\mu}}-F_{\overline{\mu}^{\prime}}\rvert_{L^{2}}\big{)}.$ (2.6)
@@
$\mathrm{Adm}(U,q):=\big{\{}x\in X\mid\text{ there exists }r>0\text{ such that for all }t\in[0,r]\text{ we have }q+tx\in U\big{\}}$ (2.7)
@@ 1. GD1
    $Dh(q;x):=\lim_{\epsilon\searrow 0}\frac{h(q+\epsilon x)-h(q)}{\epsilon}$ (2.8)
@@ 2. GD2
$D_{\mu}f(t,\mu;\nu):=\lim_{\epsilon\to 0}\frac{f(t,\mu+\epsilon\nu)-f(t,\mu)}{\epsilon}.$ (2.9)
We say that the Gateaux derivative of $f(t,\cdot)$ admits a density at the measure $\mu\in\mathcal{M}_{+}$ if there exists a continuous function $x\mapsto D_{\mu}f(t,\mu,x)$ defined on the interval $[-1,1]$ such that for all $\nu\in\mathrm{Adm}(\mathcal{M}_{+},\mu)$,
@@
We often abuse notation and identify the density $D_{\mu}f(t,\mu,\cdot)$ with the Gateaux derivative $D_{\mu}f(t,\mu)$. Implicitly identifying the continuous dual of $\mathcal{M}_{+}$ with the space of continuous functions on $[-1,1]$, this definition of Gateaux differentiability coincides with that in (2.8)-(GD2) but differs slightly from that in [18] where the density is only required to be measurable and bounded as opposed to continuous. However, all Gateaux derivative densities appearing in [18] are of the finite-volume free energy and are continuous, so this difference is insignificant.
For technical reasons, it will be convenient for the dense subset of $\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ on which we establish the Gateaux differentiability of the free energy to be composed of well-behaved measures. Recall the definition in (1.30) of the set $\mathrm{Reg}(\mathcal{M}_{+})$ of measures in $\mathcal{M}_{+}$ that admit a smooth and strictly positive density with respect to Lebesgue measure on $[-1,1]$. A key property of this set is that for any $\mu,\nu\in\mathrm{Reg}(\mathcal{M}_{+})$ and $t\in\mathbb{R}$ with $\lvert t\rvert$ small enough, we have $\mu+t\nu\in\mathcal{M}_{+}$. In particular, the set of admissible directions at $\mu\in\mathrm{Reg}(\mathcal{M}_{+})$ contains $\mathrm{Reg}(\mathcal{M}_{+})$. In this notation, the main Gateaux differentiability results for the enriched free energy proved in Sections 2.2 and 2.3 read as follows.
@@
Suppose that the sequence of enriched free energies $(\overline{F}_{N})_{N\geq 1}$ converges pointwise to some limit $f:\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}\to\mathbb{R}$ along a subsequence $(N_{k})_{k\geq 1}$. For each $t\geq 0$, if $f(t,\cdot)$ is Gateaux differentiable at $\mu\in\mathrm{Reg}(\mathcal{M}_{+})$, then $D_{\mu}\overline{F}_{N_{k}}(t,\mu,\cdot)$ converges weakly to $D_{\mu}f(t,\mu,\cdot)$.
In proving Propositions 2.1 and 2.2, it will be convenient to remember from [18, Lemmas 2.1 and 2.3] that the enriched free energy (1.22) is Gateaux differentiable at every $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ with
@@
[…]
@@
This can be first verified for functions of product form using that the Gibbs average is the conditional law of the signal $\sigma^{*}$ given the data $\mathcal{D}_{N}^{t,\mu}$, and then extended to all bounded and measurable functions by a monotone class argument as in [19, Proposition 4.1].

### v2

# arXiv:2406.15233v2[math.PR] 02 Mar 2026

@@ 1.1 Informal summary of the main results
We denote by $\mathcal{M}_{+}$ the space of non-negative measures over [1,1][-1,1] with finite total mass, and by $\mathcal{M}_{p}$ the set of probability measures with mean $\overline{m}=2p-1$,
@@
Our main results will be stated in terms of a fixed point of the functional $\Gamma_{t,\mu}:\mathcal{M}_{p}\to\mathcal{M}_{p}$, which is defined for each $t\geq 0$ and $\mu\in\mathcal{M}_{+}$ according to the following procedure. Fixing $\nu\in\mathcal{M}_{p}$, we sample $\varepsilon^{*}$ according to $P^{*}$, and conditionally on $\varepsilon^{*}$, we let $\Pi_{t,\mu}(\nu)$ be a Poisson point process with intensity measure ${(c+\Delta\varepsilon^{*}x)}\,\mathrm{d}(\mu+t\nu)(x)$; we then set
$\Gamma_{t,\mu}(\nu):=\text{Law}\Bigg(\frac{\int_{\Sigma_{1}}\varepsilon e^{-\varepsilon\overline{m}}\prod_{x\in\Pi_{t,\mu}(\nu)}(c+\Delta\varepsilon x)\,\mathrm{d}P^{*}(\varepsilon)}{\int_{\Sigma_{1}}e^{-\varepsilon\overline{m}}\prod_{x\in\Pi_{t,\mu}(\nu)}(c+\Delta\varepsilon x)\,\mathrm{d}P^{*}(\varepsilon)}\Bigg).$ (1.6)
@@
    conditionally on $(x_{i})_{i\geq 1}$, generate $\pm 1$-valued random variables $\sigma_{i}^{\ell}$ with mean $x_{i}$, independently over $i,\ell\geq 1$.
@@
A complete characterization of the limit of the mutual information has already been achieved in a number of cases. The disassortative case was settled in [2, 15] when $p=1/2$, and in [19] for arbitrary $p$. The technique in [19] is based on partial differential equations (see also [20] for a presentation of the approach), and the key point is that in the disassortative case, the non-linearity appearing in this equation is convex. We expect this complete characterization to extend to settings with more than two communities, as long as the corresponding non-linearity remains convex. This would be analogous to [11, Theorem 1.1] in the context of spin glasses, a result which itself builds upon [9, 22, 42, 43, 44, 45, 46, 47, 52]. Moreover, whenever this convexity property holds, the limit mutual information can be expressed as a variational formula. From the point of view of partial differential equations, this relates to the validity of a Hopf-Lax representation of the solution, see [18, Theorem 1.3].
@@
We expect the method presented here, primarily based on cavity calculations, to be robust to model variations such as situations involving more than two communities, and we also expect the lower bound on the mutual information obtained in [19] to generalize similarly. The results we present here and those in [19] are analogous to those obtained for spin glasses in [11] and in [39, 40] respectively. We have preferred to stick to a two-community model here to facilitate the presentation, but as an illustration of the robustness of the approach, we will at least indicate as we proceed the few small adaptations that need to be implemented in order to cover the case when then measure $P^{*}$ is arbitrary with support in [1,1][-1,1] in place of $\{-1,1\}$.
@@
In order to make further progress and obtain a full characterization of the limit mutual information in general, one must first ask what a good candidate for the limit could be. One proposal based on partial differential equations was put forward in [19]. It would of course be convenient to identify a more explicit formulation of the candidate limit. For dense versions of the problem we consider, where the average degree of the graph diverges to infinity, a sort of central limit theorem takes place at the level of each node, and one can relate the problem to a simpler setup such as that of rank-one matrix estimation in the presence of additive Gaussian noise, as shown in [17, 26] (see also [20, Section 4.5]). For the latter class of problems, a saddle-point variational formula was shown to be valid in a very wide range of settings in [14] (see also [20, Chapter 4]), building upon [4, 5, 6, 10, 13, 12, 23, 26, 27, 28, 29, 31, 32, 37, 38, 50, 51]. From the point of view of partial differential equations, this corresponds to the possibility to write a Hopf formula based on the convexity of the initial condition (as opposed to the convexity of the non-linearity for the Hopf-Lax formula, cf. [20, Theorems 3.8 and 3.13]). However, the convexity of the relevant “initial condition” in our setting is invalid [25], and as explained around [19, (1.49)], this implies that this tentative saddle-point Hopf formula is invalid in our setting with sparse graph connectivity. An alternative candidate limit for the mutual information could be considered by extending the Hopf-Lax variational formula known to be valid in the disassortative case. While we could not rule this out in the two-community case, we show here that this candidate variational formula is invalid in general. To do so, we focus on a version of the two-community problem in which the graph additionally possesses a bipartite structure. The model can be thought of as involving four communities, and it is inspired by the bipartite spin-glass model investigated in [39] (see in particular [39, Section 6]). In this context, we show that a naive generalization of the formula valid in the standard disassortative two-community case would lead to a prediction that is in contradiction with the main results of this paper.
@@ 1.2 Precise statement of the main results
We now turn to a more precise description of our main results. As in [19], our starting point is to embed the community-detection problem under consideration into a richer family parametrized by a pair $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$. The conjecture in [19] is that, up to a simple affine transformation, the limit $f(t,\mu)$ of the re-scaled mutual information between the signal and the observation solves an explicit Hamilton-Jacobi equation. As is well-known in finite dimensions, first-order Hamilton-Jacobi equations can be solved for a short time using the method of characteristics (see for instance [20, Exercise 3.10 and solution]). This method allows us to uncover the value of the solution to the equation along each characteristic line separately. At larger times though, difficulties occur because characteristic lines may intersect, and will typically not agree on what the value of the solution ought to be. Our main result can be interpreted as saying that, while this might indeed occur, there is always one characteristic line that prescribes the correct value for $f(t,\mu)$. In the language of [20, Section 3.5], one may say that the graph of $f$ always belongs to the wavefront of the equation.
@@
By this we mean that the law of $\sigma^{*}$ given the observation of $\mathcal{D}_{N}^{t}$ is described by a formula analogous to that in (1.9), with $H_{N}^{\circ}$ replaced by $H_{N}^{t}$ there. By [19, Appendix A], the difference between the free energies $\overline{F}_{N}^{\circ}$ and $\overline{F}_{N}(1)$ converges to zero as $N$ tends to infinity. The main advantage of this construction is that it allows one to study derivatives of the free energy with respect to the “time” parameter $t$. It is also convenient to study derivatives of the free energy with respect to an additional parameter denoted by $\mu\in\mathcal{M}_{+}$. Given $\mu\in\mathcal{M}_{+}$, we let $(\Lambda_{i}(N\mu))_{i\leq N}$ be independent Poisson point processes with intensity measure $N\mu$, and we let the $\mu$-dependent Hamiltonian $H_{N}^{\mu}$ on $\Sigma_{N}$ be given by
@@
For a review of Poisson point processes, we refer the reader to [20, Chapter 5]. The function $H_{N}^{\mu}$ is the Hamiltonian associated with the task of inferring the signal $\sigma^{*}$ from the data
@@
where the randomness in these two datasets is taken to be independent conditionally on $\sigma^{*}$. We recall that by [19, Appendix A], the difference between the free energy $\overline{F}_{N}^{\circ}$ in (1.11) and $\overline{F}_{N}(1,0)$ tends to zero as $N$ tends to infinity. The problem of finding the asymptotic value of the mutual information (1.7) is therefore a specific instance of the task of determining the limit of the enriched free energy (1.22).
@@
for Poisson point processes $\Pi_{\pm}(\mu)$ with respective intensity measures $(c\pm\Delta x)\,\mathrm{d}\mu(x)$ on [1,1][-1,1]. The function $\psi$ is the initial condition of the partial differential equation that was identified in [19], since it arises as the large-$N$ limit of $\overline{F}_{N}(0,\mu)$. Our first main result states that the limit of the free energy, assuming it exists, is given by the functional (1.26) evaluated at one of its critical points. As mentioned in the previous subsection, it will be possible to recast the critical-point condition as a fixed-point equation for the operator $\Gamma_{t,\mu}$ defined in (1.6), so that our first main result reads as follows.
@@
Our second result concerns the asymptotic behavior of the conditional law of the community structure $\sigma^{*}$ given the observation. We use the bracket notation $\langle\cdot\rangle$ to denote the expectation with respect to the conditional law of $\sigma^{*}$ given the observation, with $\sigma$ being the canonical random variable under $\langle\cdot\rangle$. Explicitly, this means that for every function $h:\Sigma_{N}\to\mathbb{R}$, we have
We stress that the expectation $\langle\cdot\rangle$ is itself random, since it depends on the observation. We denote by $(\sigma^{\ell})_{\ell\geq 1}=((\sigma^{\ell}_{i})_{i\leq N})_{\ell\geq 1}$ independent copies, often called replicas, of the random variable $\sigma$ under $\langle\cdot\rangle$. We also introduce the set $\mathrm{Reg}(\mathcal{M}_{+})$ of measures in $\mathcal{M}_{+}$ that admit a smooth and strictly positive density with respect to the Lebesgue measure on [1,1][-1,1],
@@
Our next result states that for most choices of $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$, we can identify a fixed point of the map $\nu\mapsto\Gamma_{t,\mu}(\nu)$ for which the equality (1.28) holds, and for which we can also identify the limit law of $(\sigma^{\ell})_{\ell\geq 1}$ after a small perturbation of the inference problem. More precisely, we will show that this can be done at any point $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathrm{Reg}(\mathcal{M}_{+})$ at which the limit free energy is Gateaux differentiable; we refer to Section 2 for a definition of Gateaux differentiability and a precise discussion of why the limit free energy is Gateaux differentiable at most points. The convergence in law of $(\sigma^{\ell}_{i})_{i\leq N,\ell\geq 1}$ is in the sense of finite-dimensional distributions. This result requires that we slightly modify the inference problem by adding a small amount of side information; the exact quantity we study then is defined in (4.6)-(4.8).
@@
Another limitation of Theorem 1.2 is that we need to add a small amount of side information to the inference problem, as described in (4.6)-(4.8), in order to identify the limit law of the posterior distribution. The perturbation is small in the sense that the amount of additional information per unknown tends to zero in the limit of large system size. Alternatively, we could consider a broader family of inference problems with side information, and then identify the limit law of the posterior distribution for most choices of the parameters defining the inference problem; see [11, Theorem 1.4] for a result in this spirit. We point out however that this perturbation is necessary for Theorem 1.2 to be valid. Indeed, even in the simpler setting of recovery of matrix tensor products, adding a small amount of side information is in some cases necessary in order to resolve certain symmetries of the model, see for instance [50, Example 2]. While it would be interesting to make further progress on this point, we are not aware of precise results even in this simpler context.
@@ Acknowledgements.
We would like to warmly thank Dmitry Panchenko and Jean Barbier for sharing their notes [7] on the free energy in the disassortative sparse stochastic block model with us, which helped us with many of the computations in Section 4. We also warmly thank Hong-Bin Chen for pointing out an error in an earlier version of the proof of Proposition 1.3.
@@ 2 Regularity properties of the free energy
In this section we discuss three regularity properties of the enriched free energy (1.22). In Section 2.1, we establish a Lipschitz continuity bound for the enriched free energy which implies in particular that the free energy restricted to the space of probability measures is Lipschitz continuous with respect to the Wasserstein distance. In Section 2.2, we prove that a Lipschitz continuous function on the space of cumulative distribution functions on [1,1][-1,1], viewed as a subset of $L^{2}[-1,1]$, is Gateaux differentiable on a dense subset of its domain. Finally, in Section 2.3, we leverage a convexity property of the free energy to prove that the sequence of derivatives of the free energy converges to the derivative of the limit free energy at any point of differentiability of said limit. We now state the three main results of this section. Since the ideas developed in these proofs will not reappear later, the reader may consider skipping these proofs on first reading.
@@
$\big\lvert\overline{F}_{N}(t,\mu)-\overline{F}_{N}(t^{\prime},\mu^{\prime})\big\rvert\leq C\big(\lvert t-t^{\prime}\rvert+\big\lvert\mu[-1,1]-\mu^{\prime}[-1,1]\big\rvert+\mu[-1,1]\lvert F_{\overline{\mu}}-F_{\overline{\mu}^{\prime}}\rvert_{L^{2}}\big),$ (2.6)
where the probability measures $\overline{\mu}$, $\overline{\mu}^{\prime}\in\Pr[-1,1]$ are such that
@@
In Sections 2.2-2.3 and throughout the paper, the notion of Gateaux differentiability that we take may differ from the most classical one, so we proceed to define it precisely. As in [11], we fix a topological vector space $X$ and a subset $U$ of $X$. We denote by $X^{*}$ the continuous dual of $X$, and we write $\langle\cdot,\cdot\rangle$ for the duality pairing. We will not necessarily take $U$ to be an open set, so, given a point $q\in U$, the set
[…]
@@ 1. GD1
    $Dh(q;x):=\lim_{\varepsilon\searrow 0}\frac{h(q+\varepsilon x)-h(q)}{\varepsilon}$ (2.9)
@@ 2. GD2
$D_{\mu}f(t,\mu;\nu):=\lim_{\varepsilon\to 0}\frac{f(t,\mu+\varepsilon\nu)-f(t,\mu)}{\varepsilon}.$ (2.10)
We say that the Gateaux derivative of $f(t,\cdot)$ admits a density at the measure $\mu\in\mathcal{M}_{+}$ if there exists a continuous function $x\mapsto D_{\mu}f(t,\mu,x)$ defined on the interval [1,1][-1,1] such that for all $\nu\in\mathrm{Adm}(\mathcal{M}_{+},\mu)$,
@@
We often abuse notation and identify the density $D_{\mu}f(t,\mu,\cdot)$ with the Gateaux derivative $D_{\mu}f(t,\mu)$. Implicitly identifying the continuous dual of $\mathcal{M}_{+}$ with the space of continuous functions on [1,1][-1,1], this definition of Gateaux differentiability coincides with that in (2.9)-(GD2) but differs slightly from that in [19] where the density is only required to be measurable and bounded as opposed to continuous. However, all Gateaux derivative densities appearing in [19] are of the finite-volume free energy and are continuous, so this difference is insignificant.
For technical reasons, it will be convenient for the dense subset of $\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ on which we establish the Gateaux differentiability of the free energy to be composed of well-behaved measures. Recall the definition in (1.30) of the set $\mathrm{Reg}(\mathcal{M}_{+})$ of measures in $\mathcal{M}_{+}$ that admit a smooth and strictly positive density with respect to Lebesgue measure on [1,1][-1,1]. A key property of this set is that for any $\mu,\nu\in\mathrm{Reg}(\mathcal{M}_{+})$ and $t\in\mathbb{R}$ with $\lvert t\rvert$ small enough, we have $\mu+t\nu\in\mathcal{M}_{+}$. In particular, the set of admissible directions at $\mu\in\mathrm{Reg}(\mathcal{M}_{+})$ contains $\mathrm{Reg}(\mathcal{M}_{+})$. In this notation, the main Gateaux differentiability results for the enriched free energy proved in Sections 2.2 and 2.3 read as follows.
@@
Suppose that the sequence of enriched free energies $(\overline{F}_{N})_{N\geq 1}$ converges pointwise to some limit $f:\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}\to\mathbb{R}$ along a subsequence $(N_{k})_{k\geq 1}$. For each $t\geq 0$, if $f(t,\cdot)$ is Gateaux differentiable at $\mu\in\mathrm{Reg}(\mathcal{M}_{+})$, then $D_{\mu}\overline{F}_{N_{k}}(t,\mu,\cdot)$ converges weakly to $D_{\mu}f(t,\mu,\cdot)$. For each $\mu\in\mathcal{M}_{+}$, if $f(\cdot,\mu)$ is differentiable at $t>0$, then $\partial_{t}\overline{F}_{N_{k}}(t,\mu)$ converges to $\partial_{t}f(t,\mu)$.
In proving Propositions 2.1 and 2.2, it will be convenient to remember from [19, Lemmas 2.1 and 2.3] that the enriched free energy (1.22) is Gateaux differentiable at every $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ with
@@
We have used $\langle\cdot\rangle$ to denote the Gibbs average associated with the enriched Hamiltonian (1.21). This means that for any bounded and measurable function $f=f(\sigma^{1},\ldots,\sigma^{n})$ of finitely many replicas,
@@
This can be first verified for functions of product form using that the Gibbs average is the conditional law of the signal $\sigma^{*}$ given the data $\mathcal{D}_{N}^{t,\mu}$, and then extended to all bounded and measurable functions by a monotone class argument as in [20, Proposition 4.1].